OpenStudy (anonymous):

Which equation is represented in the graph shown below? Ordered Pairs: 0,5 3,8 6,11 Answers: y = 5x y = x + 5 Y equals x over 5 y = x – 5

OpenStudy (phi):

same idea. Find the eq of a line using 2 points.

OpenStudy (anonymous):

easiest way is to plug in.

OpenStudy (anonymous):

5=0*5? no. 5=0+5 .yes.okay so try the other values to make sure:8=3+5? yes.11+6+5?yes. so the correct answer is y=x+5